### v4

# PhysX-3D: Physical-Grounded 3D Asset Generation

Abstract
3D modeling is moving from virtual to physical. Existing 3D generation primarily emphasizes geometries and textures while neglecting physical-grounded modeling. Consequently, despite the rapid development of 3D generative models, the synthesized 3D assets often overlook rich and important physical properties, hampering their real-world application in physical domains like simulation and embodied AI. As an initial attempt to address this challenge, we propose PhysX, an end-to-end paradigm for physical-grounded 3D asset generation. 1) To bridge the critical gap in physics-annotated 3D datasets, we present PhysXNet - the first physics-grounded 3D dataset systematically annotated across five foundational dimensions: absolute scale, material, affordance, kinematics, and function description. In particular, we devise a scalable human-in-the-loop annotation pipeline based on vision-language models, which enables efficient creation of physics-first assets from raw 3D assets. 2) Furthermore, we propose PhysXGen, a feed-forward framework for physics-grounded image-to-3D asset generation, injecting physical knowledge into the pre-trained 3D structural space. Specifically, PhysXGen employs a dual-branch architecture to explicitly model the latent correlations between 3D structures and physical properties, thereby producing 3D assets with plausible physical predictions while preserving the native geometry quality. Extensive experiments validate the superior performance and promising generalization capability of our framework. All the code, data, and models will be released to facilitate future research in generative physical AI.
Figure 1: Visualizations of our PhysXNet for phsycial 3D generation. 3D assets in our dataset have fine-grained physical property annotations, including 1) absolute scale, 2) material, 3) affordance, 4) kinematics, and 5) function descriptions (basic, functional, and kinematical descriptions).
1 Introduction
The creation of diverse and high-quality 3D assets has gained significant prominence in recent years, driven by their expanding applications across gaming, robotics, and embodied simulators. Substantial research efforts have been focused on appearance and geometry only, from high-quality 3D datasets shapenet, objaverse, objaversexl, omniobject3d, efficient 3D representations, to generative modeling. However, most of them predominantly emphasize structural characteristics while overlooking physical properties inherent to real-world objects. Given the rising demand for physical modeling, understanding, and reasoning in 3D space, we argue that a comprehensive suite for physics-grounded 3D objects is important, from upstream data annotations pipeline to downstream generative modeling.
Beyond purely structural attributes like geometry and appearance, real-world objects intrinsically possess rich physical and semantic characteristics comprising: 1) absolute scale, 2) material, 3) affordance, 4) kinematics, and 5) function descriptions. By integrating these fundamental properties with classical physical principles, we can derive critical dynamic metrics, including gravitational effects, frictional forces, contact region, motion trajectories, and interaction. However, existing datasets/annotation pipelines only offer partial solutions towards physically grounded knowledge in 3D objects that cover the entire spectrum. Recent efforts to support articulated object applications have yielded datasets like PartNet-Mobility partnetmobility, which provides 2.7K human-annotated articulated 3D models. Yet, this collection still lacks essential physical descriptors - including dimensional specifications, material composition, and functional affordances - that are crucial for physically accurate simulations and robotics applications.
To bridge this representational gap, we propose PhysXNet – the first comprehensive physical 3D dataset containing over 26K richly annotated 3D objects, as illustrated in Figure 1. Except for the object-level annotation, i.e., 1), we annotate 2) and 5) for each part. Besides, for 3), we provide the affordance rank for all parts, while we annotate the 4) detailed parameters of kinematic constraints, including motion range, motion direction, child parts, and parent parts. Besides, we introduce an extended version, PhysXNet-XL, featuring over 6 million procedurally generated and annotated 3D objects.
Most importantly, PhysXNet is built with an efficient, robust, and scalable labeling pipeline. We introduce a human-in-the-loop annotation pipeline to annotate the properties for the existing object-level 3D dataset, i.e., PartNet partnet. The pipeline proceeds in three stages: 1) target visual isolation, in which we render each component via alpha compositing to get the best visual prompts with minimized visual interference. 2) automatic VLM labeling, where a large vision‑language model (VLM) to annotate most of the properties; and 3) expert refinement, combining systematic spot‑checks with focused human annotation of complex kinematic behaviors. To the best of our knowledge, PhysXNet is the first 3D dataset with abundant properties for each part.
To bridge the modeling gap of physical-grounded 3D assets, we further introduce PhysXGen, a feedforward model for physical 3D generation. Given the fact that physical properties are spatially related to geometry and appearance, we repurpose pretrained 3D generative priors to generate physical 3D assets, enabling efficient training with reasonable generalizability. Specifically, PhysXGen leverages a dual-branch architecture to jointly model the latent correlations between 3D geometric structures and physical properties, which is naturally compatible with existing 3D native generative priors. Moreover, this formulation makes the best use of pretrained latent space, leading to plausible physical predictions while keeping the decent geometry quality from the pretrained model. Comprehensive experiments prove the promising performance of PhysXGen. We hope our work reveals new observations, challenges, and potential directions for future research in embodied AI and robotics.
To summarize, our main contributions are:
  • We pioneer the first end-to-end paradigm for physical-grounded 3D asset generation, advancing the research frontier in physical-grounded content creation and unlocking new possibilities for downstream applications in simulation.
  • We build the first physical-grounded 3D dataset, PhysXNet, and propose a human-in-the-loop annotation pipeline to convert existing geometry-focused datasets into fine-grained physics-annotated 3D datasets efficiently and robustly. In addition, we present an extended version, PhysXNet-XL, which includes over 6 million annotated 3D objects generated through procedural methods.
  • We design a dual-branch feed-forward framework, PhysXGen. It can model the latent interdependencies between structural and physical features to achieve plausible physical predictions while maintaining the native geometry quality.
2 Related Work
Table 1: Comparison of related datasets which can support research in physical 3D generation. While the ABO dataset abo contains material metadata and keywords, its object-level annotation granularity constrains part-aware applications like robotic manipulation or physical simulation. In contrast, PhysXNet provides part-level annotations.
Dataset # Objs Part anno Physical Dim Material Affordance Kinematic Description Year
ShapeNet shapenet 51K 2015
PartNet partnet 26K 2019
PartNet-Mobility partnetmobility 2.7K 2020
GAPartNet gapartnet 1.1K 2022
ABO abo 7.9K Obj-level Obj-level 2022
OmniObject3D omniobject3d 6K 2023
Objaverse objaverse 818K 2023
PhysXNet (ours) 26K Part-level Part-level 2025
PhysXNet-XL (ours) 6M Part-level Part-level 2025
2.1 3D Datasets and Benchmarks
Due to the time-consuming and expensive in realistic data collection, current large-scale 3D datasets prefer to collect data online shapenet, objaverse, objaversexl. According to the type of 3D data, existing 3D datasets can be divided into synthetic and real-world datasets. To facilitate the development of 3D vision, ShapeNet shapenet collects 51,300 CAD models. Building upon it, the PartNet dataset partnet introduces an annotation framework that provides part annotations at significantly finer granularity levels. Furthermore, PartNet-Mobility partnetmobility annotates the kinematic constraints and provides 2.7K articulated 3D objects for 3D vision, especially for embodied AI and robotics. ABO abo is a high-quality datasets with around 7.9K CAD models with fine-grained geometric and textures. Compared with prior work, it includes the physical dimension, material, and keywords. However, the material information and descriptions focus on object-level, limiting the part-aware applications. Recently, Objaverse objaverse has alleviated the scarcity of 3D data. It collects and filters over 800K 3D data. To bridge the gap between synthetic and real data, Omniobject3D omniobject3d provides over 6k high-quality 3D scans. A detailed comparison is shown in Table 1.
Despite significant advances in 3D data acquisition, prevailing 3D datasets primarily emphasize geometry and appearance fidelity or narrowly defined physical attributes, creating a critical bottleneck for developing physics-aware 3D vision models and their real-world applications. To bridge this foundational gap, we present PhysXNet – a 3D dataset with comprehensive physical properties encompassing physical dimension, part-level material, affordance rank, kinematic parameters, and part-level description. Furthermore, we extend our dataset with PhysXNet-XL, comprising more than 6 million annotated 3D objects created via procedural generation.
2.2 3D Generative Models
As one of the most representative optimization-based method in 3D generation, DreamFusion poole2022dreamfusion proposed the SDS loss function. By utilizing the prior knowledge of the 2D diffusion model, it achieves impressive generative performance. Despite various works, optimization-based methods still suffer from the multi-face Janus problem and low optimization efficiency. Recently, benefiting from its impressive efficiency and robustness, feed-foward models trellis, lgm, instantmesh, zlabor, 3dtopia, 3dtopiaxl, difftf, difftf++ have gained more and more attention. However, those methods still focus on geometry and appearance quality, neglecting the physical properties of 3D assets.
2.3 Articulated and Physical 3D Object Modeling
Articulated object modeling mainly consists of tasks like perception, reconstruction, and generation. Some works try to estimate articulation pose liu2022toward and identify articulation parts zengvisual, while others sun2024opdmulti focus on learn joint parameters from images. In the reconstruction field, existing works try to reconstruct articulated models from RGB chen2024urdformer, RGBD weng2024neural, and point cloud hsu2023ditto. Recently, some methods have tried to generate articulated 3D assets by utilizing a vision-language model le2024articulate, cao2025physx or adopting an optimization-based framework qiu2025articulate. To bridge the critical gap between existing methods with real applications, many works aim to incorporate the physical properties into 3D modeling. Some works try to learn material parameters from videos zhong2024reconstruction or images zhai2024physical, while other methods aim to introduce physical guidance via simulation mezghanni2022physical, xie2024physgaussian or physical principles guo2024physically.
In contrast to fragmented paradigms in physical 3D modeling, this work introduces PhysXGen – a unified physics-integrated generative framework capable of learning cross-property consistency to generate 3D assets with all necessary physical properties. By exploiting the relationship between physical and structural features, our method achieves promising performance in physical 3D generation.
3 PhysXNet Dataset
In this section, we will introduce physical properties and the human-in-the-loop annotation pipeline. Besides, we will report the statistics and distribution of PhysXNetand PhysXNet-XL.
Figure 2: Top: Definition of properties in PhysXNet . By defining and annotating properties across three categories, common physical quantities can be systematically calculated to enable physical simulations. Bottom: Overview of our human-in-the-loop annotation pipeline. We utilize GPT-4o to gather foundational raw data, which is subsequently verified through human oversight. The kinematic parameters are then rigorously determined and finalized through human review.
3.1 Definition of Physical Properties
As shown in Figure 2, we systematically categorize object properties into three progressive stages: a) Identification - determining the basic nature of the object; b) Function - understanding its potential applications; and c) Operation - detailed usage methodologies. To streamline the annotation process, we posit that the internal composition of a component is homogeneous, exhibiting uniform property invariance throughout its structure. For stages a), we set absolute scaling and material (material name, Young’s modulus, Poisson’s ratio, and density). Besides, for b), we establish functional affordance analysis and function descriptions (basic, functional, and kinematic descriptions). Finally, we use kinematic parameter quantification to represent c). Specifically, we grade the priority of being touched on all available parts to obtain the affordance score for all parts from 1 to 10. We set five possible kinematic types: A. No movement constraints (like water in a bottle), B. Prismatic joints (like a drawer), C. Revolute joints (like a laptop), D. Hinge joint (like a hose in a shower system), or E. Rigid joint and a combined kinematic type: CB. Revolute and Prismatic joints (like a lid of a bottle). Except for A and E, we will annotate the parent, child parts, and detailed kinematic parameters (such as rotation direction, rotation range, and so on). Note that, due to the challenges in precisely quantifying the absolute physical movement range of B, we use the movement range within the 3D coordinate system. Besides, to avoid the unnecessary and meaningless annotation of over-fine-grained parts in PartNet, we merge the tiny parts whose vertices and area are smaller than a pre-defined threshold with their neighboring parts. We manually refine the results of the merging process to ensure that the merged outputs are reasonable and consistent.
Figure 3: Statistics and distribution of PhysXNet and PhysXNet-XL. (a) Distribution histogram of part number in PhysXNet. (b) Dimensional distribution analysis in PhysXNet, showing physical measurements (length/width/height) frequency. (c) Proportional composition of kinematic states and material, including density, Young’s modulus, and Poisson’s ratio distribution in PhysXNet, visualized through sectoral ratios. (d) Tag frequency statistics for prevalent object labels in PhysXNet-XL. (e) Component-Category distribution of procedurally generated 3D objects in PhysXNet-XL.
3.2 Human-in-the-loop Annotation Pipeline
Following the establishment of target annotation specifications, we implement a systematic and streamlined semi-automated annotation framework, structured into two distinct operational phases (see Figure 2): 1) Preliminary Data Acquisition and 2) Kinematic Parameter Determination. Specifically, we utilize GPT-4o to obtain the basic information. Besides, to ensure the quality of raw data, a human candidate will check and refine the output of the vision-language model (VLM).
For the second phrase, we split it into four subtasks: (2.a) calculate contact region, (2.b) plane fitting, (2.c) candidate generation and selection, and (2.d) kinematic parameters. Note that (2.c) and (2.d) are accomplished by human candidate. For all constraint movable parts (kinematic type is not A or E), we will calculate the contact region with the neighboring parts. We first extract point cloud data from the child-parent mesh pair, formally designated as $P_{c}$ and $P_{p}$, respectively. The workflow subsequently calculates Euclidean distance between points in $P_{c}$ and $P_{p}$, followed by spatial filtration that eliminates point pairs failing to meet a predetermined distance threshold. Subsequently, we employ a plane-fitting algorithm. We sample several axes uniformly on the fitted plane as candidates. Note that for kinematic type C, we additionally need to determine the location of the rotation axis. Therefore, we will perform a k-means algorithm in the contact region for type C to generate several candidates. After selecting the candidate location, we can finalize the kinematic parameters.
3.3 Statistics and Distribution of PhysXNet
Comprises over 26K physical 3D objects, the part number of objects in PhysXNet exhibits a long-tailed distribution illustrated in Figure 3, where each object contains an average of around 5 constituent parts. Besides, we document the length-width-height distributions of objects in (b). Given that PhysXNet encompasses objects spanning from relatively small-scale indoor entities to large-scale outdoor structures, the physical dimension exhibits significant variation among objects. For kinematic types and material in PhysXNet, we show detailed proportional composition. Note the density in our PhysXNetadheres to the metric standardization framework, i.e., $g/cm^{3}$. Furthermore, Figure 3 (d) shows the frequency of the popular object tags, including the name and category. Finally, we also report the component category in our procedurally generated 3D objects, including a) intra-category combination: cabinet, bottle, faucet, chair, oven, shower, knife, table, and laptop; b) cross-category combination: drawer and door. More details about PhysXNet-XL are released in the appendix.
4 PhysXGen Framework
As mentioned above, physical 3D generation is still a challenging and promising task. Most prior works only focus on a single or specific physical property. In this section, we aim to build a unified generative framework to generate physical 3D assets directly. While our PhysXNet dataset contains 26K assets, this scale remains insufficient for training SOTA generative architectures from scratch. Therefore, we leverage a model pre-trained on massive geometry-only 3D scans and fine-tune it to adapt to physical 3D generation. Building upon the well-established 3D representation space of it, we present PhysXGen, a novel yet straightforward framework that combines physical properties with geometry and appearance shown in Figure 4. Our approach achieves this dual objective by simultaneously integrating fundamental physical properties into the generation process while optimizing the structural branch through targeted fine-tuning. This joint optimization enables the production of physically consistent 3D assets that maintain impressive geometry and appearance fidelity.
Figure 4: The architecture of PhysXGen framework. PhysXGen features a two-stage architecture comprising: a physical 3D VAE framework for latent space learning, and a physics-aware generative process for structured latent. The former focuses on establishing a compressed yet information-rich latent representation that encodes physical properties, while the latter specializes in generating physical latents.
4.1 Physical 3D VAE Encoding and Decoding
In this subsection, we take the textured mesh output as an example. To reduce the influence caused by the domain gap between geometric and physical latent space, we build a similar physical VAE for property encoding, following trellis. Besides, considering the interdependencies among physical properties, we encode them into a unified latent space. We adopt 4 physical properties: physical scaling (converted by physical dimension) $P_{dim}\in\mathbb{R}^{N\times 1}$, affordance priority $P_{aff}\in\mathbb{R}^{N\times 1}$, density $P_{\rho}\in\mathbb{R}^{N\times 1}$, and kinematic parameters $P_{mov}\in\mathbb{R}^{N\times 11}$ (including child $\mathbb{R}^{N\times 1}$ and parent group index $\mathbb{R}^{N\times 1}$, movement direction $\mathbb{R}^{N\times 3}$, movement location $\mathbb{R}^{N\times 3}$, movement range $\mathbb{R}^{N\times 2}$, and kinematic type $\mathbb{R}^{N\times 1}$), where N is the number of voxel. The physical properties ($P_{phy}\in\mathbb{R}^{N\times 14}$) can be obtained by channel-wise concatenation. For the function descriptions, we adopt the CLIP model radford2021learning to obtain the text embedding. Similarly, the description features ($P_{sem}\in\mathbb{R}^{N\times 768\times 3}$) are formed by concatenating the basic, functional, and kinematic description embeddings. Besides, the structural branch adopts the DINOv2 to extract features. Therefore, the dimensions of structural feature is $P_{aes}\in\mathbb{R}^{N\times 1024}$. For clarification, we denote the pretrain VAE encoder and decoder as $\mathcal{E}_{aes}$ and $\mathcal{D}_{aes}$ while the physical VAE encoder and decoder as $\mathcal{E}_{phy}$ and $\mathcal{D}_{phy}$. The physical latent $P_{plat}\in\mathcal{R}^{N\times 8}$ and structured latent $P_{slat}\in\mathcal{R}^{N\times 8}$ can be formulated as follows:
$\begin{split}P_{plat}=\mathcal{E}_{phy}(P_{phy},P_{sem}),~P_{slat}=\mathcal{E}_{aes}(P_{aes})\end{split}~.$ (1)
To study the effects of physical properties on geometry and appearance quality, we introduce a branch from $\mathcal{D}_{phy}$ to $\mathcal{D}_{aes}$ via a residual connection. We will analyze the performance of the independent and dependent VAE decoder in the experiments. After decoding the structured and physical latents, we can implement a loss function $\mathcal{L}$ as follows:
$\begin{split}\mathcal{L}_{vae}=\mathcal{L}_{aes}^{color}+\mathcal{L}_{aes}^{geometry}+\mathcal{L}_{phy}+\mathcal{L}_{sem}+\mathcal{L}_{kl}+\mathcal{L}_{reg}\end{split}~,$ (2)
where $\mathcal{L}_{aes}^{color}$ and $\mathcal{L}_{aes}^{geometry}$ represent the color loss (including L2loss, lpip loss) and geometry loss (including mask, normal, and depth loss). For $\mathcal{L}_{phy}$ and $\mathcal{L}_{sem}$, we normalize the groundtruth respectively and adopt a L2 loss. $\mathcal{L}_{kl}$ aims to constraint the distribution of $P_{plat}$ while $\mathcal{L}_{reg}$ can reduce the unnecessary structures of textured mesh.
4.2 Physical Latent Generation
Following the acquisition of the compressed physical latent representation, we construct a transformer-architecture diffusion model to jointly generate physical and structural attributes. To effectively leverage the inherent correlations between physical properties and structural features while maintaining compatibility with pre-trained components, we implement a dual-branch architecture that integrates structural guidance through residual connections. Specifically, the additional branch from the structural module is fused with the primary physical generation module via learnable skip-connection layers, enabling cross-domain feature interaction. Comprehensive ablation studies quantitatively validate the design rationale through systematic component comparisons. Following trellis, we adopt the Conditional Flow Matching (CFM) as the objective of optimization. Therefore, the loss of the geometric branch is formulated:
[…]
where $\epsilon$ and $t$ represent the noise and timestep while $x_{0}$ is sampled from $P_{slat}$. Adopting a similar objective for the physical branch, the final loss of the latent diffusion model can be calculated as: $\mathcal{L}_{diff}=\mathcal{L}_{aes}+\mathcal{L}_{phy}$.
5 Experiments
5.1 Implementation details
In our experiments, we partition PhysXNet dataset into 24K training samples, 1K validation samples, and 1K test cases. By analyzing the performance on the test cases, we can evaluate the generalizability of our method. During the VAE and diffusion model training, we adopt AdamW with an initial learning rate of $1\times 10^{-4}$ to optimize the models. The inherent correlation between geometric configuration and physical properties in our methodology creates a critical dependency where the structural fidelity of the 3D representation will affect the final generative performance. In this paper, we repurpose the geometry- and appearance-rich structural space of TRELLIS trellis for our task. Our PhysXGen is trained on 8 NVIDIA A100 GPUS. More details about the architecture are released in the supplementary.
5.2 Evaluation Metrics
Physical properties evaluation. Our framework establishes a multi-property feature space encompassing five core attributes: absolute scale, material, affordance, kinematics, and function descriptors. Note that the kinematics attribute manifests as dual configuration parameters: 1) structural grouping (parent-child part hierarchies) and 2) kinematic parameters. Specifically, we evaluate absolute scale using Euclidean distance, density and affordance images via Peak Signal-to-Noise Ratio (PSNR), kinematics with instantiation distance lei2023nap, and functional description through PSNR on cosine similarity score maps.
Table 2: Quantitative comparison of different methods on the test sets of our PhysXNet. There are two types of evaluations: structural and physical property evaluations. PhysPre represents a separate physical property predictor after TRELLIS.
Methods Geometry Absolute scale $\downarrow$ Material $\uparrow$ Affordance $\uparrow$ Kinematic parameters Description $\uparrow$
PSNR $\uparrow$ CD $\downarrow$ F-Score $\uparrow$     COV $\uparrow$   MMD $\downarrow$
TRELLIS trellis 24.31 13.2 76.9
TRELLIS + PhysPre 24.31 13.2 76.9 13.21 8.63 7.23 0.24 0.12 6.55
PhysXGen 24.53 12.7 77.3 7.24 13.01 11.30 0.33 0.08 10.11
Figure 5: Visualization of the generated results. Given a single image as the prompt, our PhysXGen can generate the physical-grounded 3D assets.
Geometry evaluation. For appearance evaluation, we sample 30 random views from a unit sphere to calculate the mean PSNR. Besides, to evaluate the quality of geometry, we calculate the standard shape metrics of Chamfer Distance (CD) ($\times 10^{-3}$) and F-score (FS) ($\times 10^{-2}$) with thresholds of 0.05.
Table 3: Ablation studies about the physical 3D VAE and diffusion model. Dep-VAE and Dep-Diff represent the model that utilizes the interdependencies between structural and physical information. Thus, Trellis+PhysPre and PhysXGen are corresponding to the first and last lines.
Dep-VAE Dep-Diff Geometry Absolute scale $\downarrow$ Material $\uparrow$ Affordance $\uparrow$ Kinematic parameters Description $\uparrow$
PSNR $\uparrow$ CD $\downarrow$ F-Score $\uparrow$     COV $\uparrow$   MMD $\downarrow$
24.31 13.2 76.9 13.21 8.63 7.23 0.24 0.12 6.55
24.31 13.2 76.9 12.01 10.69 8.95 0.26 0.11 7.71
24.32 12.9 77.0 10.57 9.86 9.32 0.28 0.11 7.54
24.53 12.7 77.3 7.24 13.01 11.30 0.33 0.08 10.11
Figure 6: Qualitative comparison of different methods. Compared with our baseline, PhysXGen achieves significant improvements, clearly demonstrating its strong performance in physics-grounded 3D generation.
5.3 Quantitative Results
As shown in Table 2, we implement the quantitative evaluations on two types of metrics: 1) geometry and appearance evaluation; and 2) physical properties evaluation. Note that TRELLIS+PhysPre is our baseline that adopts the independent structure to predict the properties. Compared with the separate physical property predictor, our PhysXNet utilizes the correlation between physical and pre-defined 3D structural space, achieving significant improvement in physical property generation while enhancing the aesthetic quality.
Ablation studies.  The core design of our framework is to integrate both geometry and physics in 3D modeling. Therefore, we conduct ablation studies to validate its effectiveness (reported in Table 3). By introducing geometry and appearance features in the diffusion model, the generative model can gain improvement in physics generation compared with the independent models, PhysPre. Additionally, the correlation between geometry and physics in VAE can enhance the geometry of generated assets. Finally, relying on the dual-architecture and joint training, our PhysXGen obtains impressive performance in all physical property generation.
5.4 Qualitative Results
Figure 5 showcases the physical-grounded 3D assets generated by our PhysXGen. By learning the interdependencies between physical and structural space, PhysXGen achieves impressive performance in generating physical properties. Besides, we perform qualitative comparisons with our baseline shown in Figure 6. As we mentioned above, for absolute scaling, we use the Euclidean distance while we adopt PSNR to evaluate the material maps, affordance maps, function description similarity score maps. By utilizing the interdependencies between physical properties and structural information, especially geometry, our PhysXNet obtains higher overall scores. Furthermore, our PhysXGen can distinguish the properties of different parts and achieve more stable and robust performance in physical property generation of neighboring structures, especially in function description, material , and affordance. More experimental results are shown in the supplementary.
6 Conclusion
In this paper, to fill the gap between existing synthesized 3D assets and real-world applications, we propose an end-to-end generative paradigm for physical-grounded 3D asset generation, including the first physical-grounded 3D dataset and the novel physical property generator. Specifically, we develop a human-in-the-loop annotation pipeline that transforms current 3D repositories into physics-enabled datasets. Meanwhile, the novel end-to-end generative framework, PhysXGen, can integrate physical priors into structural-focused architectures to achieve robust generation performance. Through comprehensive experiments on PhysXNet, we reveal the fundamental challenges and direction in physical 3D generation. We believe that our dataset will attract research attention from different communities, including but not limited to embedded AI, robotics, and 3D vision.
Limitations and Future works. Despite impressive performance, our method exhibits limitations in learning fine-grained properties and suffers from artifacts. In our future work, we will try to handle it. Besides, we will include more 3D data from synthetic to real to improve the diversity of our dataset and integrate additional physical properties and kinematic types to better simulate material behavior and movement.
Acknowledgment
This study is supported by the National Key R&D Program of China (2022ZD0160201), and Shanghai Artificial Intelligence Laboratory. This study is also supported by the Ministry of Education, Singapore, under its MOE AcRF Tier 2 (MOET2EP20221-0012, MOE-T2EP20223-0002). This research is also supported by cash and in-kind funding from NTU S-Lab and industry partner(s).
7 Implementation details
Structure of PhysXGen. In this section, we present the architectural details and implementation specifics of PhysXGen. To maintain consistency with the established pre-trained geometry space, our geometrical decoder preserves the original hyperparameter configuration from  trellis, ensuring effective utilization of pre-trained weights. For the physical processing components, we implement structurally symmetric encoder-decoder pairs (as detailed in Table 4). Notably, our physical generator employs a streamlined transformer architecture with 14 processing blocks rather than the conventional 24-block configuration to achieve satisfactory performance with lower computational overhead.
Texture retrieval for PhysXGen training. While the 3D objects in PartNet partnet inherently lack surface texturing data, we retrieve compatible UV texture coordinates from ShapeNet shapenet. For instances where no corresponding texture exists in ShapeNet, we employ the grey color as their texture information.
Figure 7: Qualitative comparison of different annotation setting.
Table 4: Hyper-parameters of main modules.
Model resolution model channel latent channels num_blocks num_heads mlp_ratio window_size
Geometry decoder 64 768 8 12 12 4 8
Physics decoder 64 2048 8 4 16 4 8
Physical encoder 64 768 8 4 12 4 8
Table 5: Quantitative comparison against GPT-based method.
Methods Geometry Absolute scale $\downarrow$ Material $\uparrow$ Affordance $\uparrow$ Kinematic parameters Description $\uparrow$
PSNR $\uparrow$ CD $\downarrow$ F-Score $\uparrow$     COV $\uparrow$   MMD $\downarrow$
TRE trellis + PartField liu2025partfieldlearning3dfeature + GPT 24.31 13.2 76.9 8.81 7.95 6.73 0.09 0.24 14.31
PhysXGen 24.53 12.7 77.3 7.24 13.01 11.30 0.33 0.08 10.11
8 Details of human-in-the-loop annotation pipeline
In this section, we detail the technical configuration of our 3D annotation pipeline. For part-aware geometric annotation, two distinct methodologies were evaluated: segmentation-based and part-based approaches. The segmentation-based method employs multi-view projective rendering to establish inter-part spatial relationships in 2D projections. To avoid the occlusion caused by the number label in the rendered images, we input the index image for reference. While effective for macro-structural analysis, this approach demonstrates limitations in capturing occluded components and accurately resolving fine geometric details that fall below the effective pixel resolution threshold. Conversely, the part-based paradigm demonstrates superior robustness in occluded and micro-part annotations. However, this methodology introduces scalability challenges when processing complex assemblies with high part counts, as it requires rendering a separate image for each individual component - a process that incurs increasing expense with increasing part number.
To avoid the expensive annotation of part-based annotation and build a robust and efficient annotation framework, we implement the following preprocessing pipeline: First, we normalize the 3D object’s spatial coordinates to the [-1, 1] range through proportional scaling and translation. Subsequently, we perform geometric simplification by filtering and merging insignificant components based on dual criteria: surface fragments with area $\leq$ 0.2, or those simultaneously satisfying face count $\leq$ 100 and area $\leq$ 0.06, are systematically merged with their topologically adjacent regions. After removing the unnecessary parts, we perform the part-based annotation. Figure 7 shows the qualitative comparison of two annotation paradigms. The segmentation-based annotation pipeline exhibits a higher propensity for generating inconsistent structural interpretations. A representative case involves Part 9, which demonstrates translational movement relative to Part 2 (annotation B) instead of maintaining the expected rigid connection (annotation E). Besides, Part 6 can move relative to the base of the drawer (Part 2, Part 10, Part 13, or Part 14) rather than Part 9. Finally, we adopt the part-based annotation pipeline due to its robustness. Furthermore, we show the system prompt for part-based annotation (see Listing 1). By annotating from global to local, we can get better annotations.
Figure 8: Workflow of our procedural generation method. Leveraging procedural generation within PhysXNet, we automatically generate over 6 million physically plausible 3D assets, forming an extended dataset denoted as PhysXNet-XL.
9 Procedural generation in PhysXNet-XL
To facilitate robust and diverse physical 3D generation, we devise a set of procedural generation rules aimed at synthesizing a broad spectrum of physically plausible 3D assets. These rules are categorized into two types: a) intra-category procedural generation and b) cross-category procedural generation. To ensure the performance of procedural generation, we choose the parts that typically exhibit similar physical properties. For a), we target object classes with structural variability, including cabinets, tables, bottles, faucets, chairs, ovens, showers, knives, and laptops. For b), we identify drawers and doors as modular components that can be flexibly integrated into different object types to enhance compositional diversity. Figure  8 shows the workflow of our procedural generation method. Specifically, we identify the connected regions between the original object and the target part. To ensure structural and physical consistency, we adapt the scale of the new component to align it appropriately with the geometry of the base structure. Finally, there are more than 6 million physical 3D objects in our PhysXNet-XL. We will try to extend more categories in our future work.
10 More experimental results
10.1 Comparison with GPT-based baseline
To evaluate the capabilities of our proposed method, PhysXGen, in generating physically-grounded 3D assets, we conduct comprehensive qualitative and quantitative comparisons against a GPT-based baseline pipeline comprising Trellis trellis, PartField liu2025partfieldlearning3dfeature, and GPT-4o. Under this benchmark framework, given an input image prompt, Trellis first generates textured 3D meshes with complete geometric and appearance representations. These assets are subsequently processed by PartField to perform fine-grained part segmentation, followed by a GPT-based physical property assignment module that predicts material parameters and dynamic attributes for each identified part. As shown in Table 5, our method exceeds the GPT-based method in geometry and most physics metrics. Across the four evaluation dimensions of absolute scale, material, kinematics, and affordance, PhysXGen demonstrates significant performance gains over the GPT-based baseline, achieving relative improvements of 24%, 64%, 28%, and 72%, respectively. In function description, our PhysXGen performs reduced robustness compared to GPT-4o, primarily attributable to its training on a relatively smaller dataset, i.e., PhysXNet. Furthermore, we visualize the generated results of the
Figure 9: Qualitative comparison of different methods. Compared with the existing method, our method achieves robust performance in generating physical 3D assets.
Figure 10: Qualitative comparison of different architectures.
GPT-based baseline and our PhysXGen in Fig. 9. The qualitative evaluations prove the impressive performance in physical-grounded 3D asset generation, especially in kinematics and affordance.
10.2 Qualitative comparison among different architectures
Additionally, we implement the qualitative evaluations of the different architectures in ablation studies (see Figure 10). By integrating the correlation between geometry and physics in VAE and diffusion model, the generative performance of physical properties is improved gradually. In material, kinematics, and affordance, our PhysXGen is more stable and accurate in determining the target region with fewer artifacts.
Figure 11: Error distribution of different physical properties.
11 Further analysis on challenges in physical-grounded 3D generation
In this section, we analyze the new challenges in physical-grounded 3D asset generation. For clarification, we summarize the special challenges in physical property generation.
Absolute scale: Our experimental results with PhysXGen reveal a constraint in absolute scale prediction: conventional normalization strategies prove inadequate for handling the inherent challenges of dimensional distribution. The absolute scale measurements exhibit a long-tailed distribution spanning three orders of magnitude (1-1000 cm), with a concentration of most samples below 300 cm. This long-tailed distribution makes linear normalization suboptimal due to its poor preservation of relative scale differences in the predominant sub-300 cm range. While logarithmic normalization presents a compelling alternative for handling its span, direct implementation would disproportionately compress the feature space where most of the objects reside (1-300 cm range), potentially diminishing discriminative power within this critical operational regime. Figure 11 shows the error distribution in Absolute scale. Our PhysXGen is hard to maintain robustness in generating extremely large objects.
Material and Affordance: Our analysis further identifies analogous normalization challenges in material density prediction (0-10 g/cm³ range), though with diminished urgency compared to absolute scaling due to the constrained parameter space. However, a more critical limitation emerges in physical property coherence: both affordance estimation and material prediction exhibit spatial inconsistency artifacts as shown in Fig. 10. Besides, we report the error distribution in the two metrics in Fig. 11. As evidenced by the distribution figure, artifact-induced perturbations manifest as spatially scattered data points in the distribution. Furthermore, although morphological post-processing can enhance the generated results, the inconsistency in the physics space of neighboring regions may obstruct further improvement in physical-grounded 3D asset generation.
Kinematics: As the fine-grained physical properties, we split the kinematics into several parameters: 1) child part; 2) parent part; 3) movement type: A. No movement constraints (like water in a bottle), B. Prismatic joints (like a drawer), C. Revolute joints (like a laptop), D. Hinge joint (like a hose in a shower system), or E. Rigid joint; 4) kinematic parameters including rotation/movement direction, location of rotation axis, rotation/movement range. For challenges 1) and 2), the inherent difficulty in determining the number of parts during generation precludes effective implementation of classification-based loss. Consequently, in our method, our adoption of regression-based prediction inadvertently introduces artifacts in hierarchical part determination (parent-child relationships). More critically, the absence of explicit mapping between 3D coordinate systems and geometric structural features increases the difficulty in building kinematics space and inserting the correlation between physics and geometry shown in Fig 11.
Function description: Our framework leverages CLIP radford2021learning for text embedding extraction, subsequently performing dimensionality reduction through 3D VAE to establish a compressed latent space. This architecture enables joint learning of all physical properties. However, the non-invertible nature of CLIP’s encoder-only architecture fundamentally restricts embedding-to-prompt disentanglement, thereby constraining interpretability in downstream 3D semantic reasoning tasks. Meanwhile, compared with other physical properties, text embedding is more complex to learn and generate. As shown in Fig 11, the error in normalized function descriptions is larger than other properties. Furthermore, while encoder-decoder foundation models like T5 raffel2020exploring theoretically permit decoding, their high-dimensional embedding spaces impose prohibitively expensive computational overhead for cross-domain alignment with physical properties.
12 Ethics statement
Human Annotator Ethical Concerns: All annotations used in this study were performed by the authors. No external participants were involved, and no personal or sensitive data were collected. According to our institution’s ethical guidelines, this study does not constitute human subjects research and therefore did not require IRB approval.
Clarification of dataset license: Since 3D data in PhysXNet and PhysXNet-XL are derived and modified from PartNet and ShapeNet, users are required to comply with the ShapeNet license terms [^1] https://shapenet.org/terms.
Potential biases: While our dataset, PhysXNet, offers a new collection of 3D objects annotated with rich physical properties, we acknowledge that the dataset may contain representational biases, which we plan to address in future work. Moreover, we caution against the direct application of our methods in safety-critical domains (e.g., autonomous robotics or medical devices) without rigorous validation, as errors or inaccuracies in physical property predictions could result in adverse outcomes. Finally, since part of the annotation process involves vision-language models (VLMs), specifically GPT-4o (costs over \$1k), the dataset annotations may reflect biases inherent in these models despite subsequent human verification.
Listing 1: System prompt for part-based annotation (GPT)
You have a good understanding of the structure of an articulated object. Your job is to assist the user in analyzing the properties of it. Specifically, the user will give you images of parts, and your task is to recognize the articulated object and analyze the parts of that object. You should find a similar physical 3D object in the real world. Based on human knowledge of it, you should give your answer about the information as follows:
Object-level:
(1) name, category, and dimension (length*width*height, in cm) of the articulated object.
Part-level:
Part_1 (image_1):
(1) Label, name, material, density (g/cm^3) of the part.
(2) priority rank of being touched when using this object based on human preference.
(3) labels of all neighboring parts.
(3.1) assign a movement type for each group between Part_1 and its neighboring parts (A. merely touch and no movement constraints, B. relative translationally move, C. rotation about an axis, D. rotation about a point, or E. rigid constraint). If the movement type is B, C, or D, output the parent and child parts.
(3.2) assign a movement type for each group between Part_1 and its neighboring parts (A. merely touch and no movement constraints, B. relative translationally move, C. rotation about an axis, D. rotation about a point, or E. rigid constraint). If the movement type is B, C, or D, output the parent and child parts.
(4) summarize the basic information (including material, physical dimension, category, and name), functional, movement description, and priority of being grasped description.
Part_2 (image_2):
(1) Label, name, material, density (g/cm^3) of the part.
(2) priority rank of being touched when using this object based on human preference.
(3) labels of all neighboring parts.
(3.1) assign a movement type for each group between Part_2 and its neighboring parts (A. merely touch and no movement constraints, B. relative translationally move, C. rotation about an axis, D. rotation about a point, or E. rigid constraint). If the movement type is B, C, or D, output the parent and child parts.
(3.2) assign a movement type for each group between Part_2 and its neighboring parts (A. merely touch and no movement constraints, B. relative translationally move, C. rotation about an axis, D. rotation about a point, or E. rigid constraint). If the movement type is B, C, or D, output the parent and child parts.
(4) summarize the basic information (including material, physical dimension, category, and name), functional, movement description, and priority of being grasped description.
For example:
"object_name": "Rifle",
"category": "ToyGun",
"dimension": "80*10*25",
"parts": [
"label": 1,
"material": "Plastic",
"density": "1.2 g/cm^3",
"name": "Foregrip",
"priority_rank": 2,
"neighbors": [
"labels_of_movement_group": "1-8",
"movement_type": "E",
"Basic_description": "Its a foregrip of a Rifle made of plastic.",
"Functional_description": "It can control the ...",
"Movement_description": "It cannot move normally...",
"Grasped_description": "Most likely to be grasped or handled.",
"label": 2,
"material": "Plastic",
"density": "1.2 g/cm^3",
"name": "Stock",
"priority_rank": 5,
"neighbors": [
"labels_of_movement_group": "2-8",
"movement_type": "B",
"parent_label": 8,
"child_label": 2
"Basic_description": "Its a foregrip of a Rifle classified as a gun. It is a big part of the object made of plastic.",
"Functional_description": "It can be grasped to control the object...",
"Movement_description": "It cannot move normally...",
"Grasped_description": "Less likely to be grasped.",
Remember:
(1) Do not answer anything not asked.
(2) You should base on the physical 3D object in the real world to analyze the properties and movement of the object.
(3) You should purely based on its function to detremine the movement type of parts.
(4) You should prefer to analyze the rendered object as a real 3D object rather than a toy model.
(5) You should assign the priority rank of being grasped from 1 to 10. The most likely part to be touched is 1.
(6) You should consider the function rather than the area or name of the target part to determine the priority rank of being grasped.
(7) The target part uses red color while the other parts use grey color.
(8) You should output full JSON including all parts.